# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 5.1. Approximating ${\mathcal{Z}}_{N}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$
To bound (5.10), a fundamental tool is still the renewal property (5.8). For the cases $\mathrm{i}_{1}<K_{\varepsilon}$ and $\mathrm{i}_{k}>\frac{1}{\varepsilon}-K_{\varepsilon}$, since each ${\mathcal{T}}_{\varepsilon}(\mathrm{i})$ is of length $\varepsilon$, we have an integral on $[0,K_{\varepsilon}\varepsilon]$, whose length tends to 00 by the choice of $K_{\varepsilon}$ in (2.38). For the case that there exists consecutive $\mathrm{i}-\mathrm{i}^{\prime}<K_{\varepsilon}$ and $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}<K_{\varepsilon}$, we need to consider separately the cases (1) $\mathrm{i}^{\prime}-\mathrm{i}\geq 2$ and $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}\geq 2$; (2) $\mathrm{i}^{\prime}-\mathrm{i}=1$ or $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}=1$. The case (1) is relatively simple, since we can bound the transition kernels between mesoscopic intervals from below uniformly. The case (2) needs a delicate computation, the readers are recommended to refer to [CSZ23, Section 5.1] for details. ∎
5.2. Approximating ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}$ and $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
@@
(5.11) $\limsup\limits_{N\to\infty}$ $\left\|\left({\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}-{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}\right)(\varphi,\psi)\right\|_{2}^{2}\leq C\Big{(}\frac{1}{\log\frac{1}{\varepsilon}}+\frac{(\log\frac{1}{\varepsilon})^{2}}{K_{\varepsilon}^{2}}\Big{)}\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2},$
(5.12) $\lim\limits_{N\to\infty}$ $\left\|\left({\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}-\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)\right)(\varphi,\psi)\right\|_{2}^{2}=0.$
@@
Step 1. To expand ${\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$, we bind $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ together via a transition probability $p_{d^{\prime}-f}(0)$, where $d,f\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i})$ and $d^{\prime},f^{\prime}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}^{\prime})$ with $\mathrm{i}^{\prime}-\mathrm{i}<K_{\varepsilon}$. Since each i=($\vec{\mathrm{i}}=(\mathrm{i},\mathrm{i}^{\prime})$ can at most refer to 2 mesoscopic intervals. By expanding ${\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$, the terms refer to indices ${\mathbf{I}}=(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$ with $(\log\frac{1}{\varepsilon})^{2}\leq k\leq 2(\log\frac{1}{\varepsilon})^{2}$ will be added, and the error has been controlled by (5.3) and (5.10).
Step 2. Then we replace the transition kernels by the heat kernels. Note that each time we bind two $X_{d_{j},f_{j}}$ and $X_{d^{\prime}_{j},f^{\prime}_{j}}$ together to get a $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})$, we have an extra $\sqrt{\varepsilon N}$. For the kernel connecting $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j-1})$ and $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})$, we have an approximation $\sqrt{\varepsilon N}p_{d_{j}-f^{\prime}_{j-1}}(0)\approx g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)$. We need to show that such kernel replacement is controllable. By [LL10, Theorem 2.3.5], since the random walk has finite third moment, it follows that for fixed $\varepsilon>0$ and $N$ large enough,
$\Big{|}p_{d_{j}-f^{\prime}_{j-1}}(0)-\frac{1}{\sqrt{\varepsilon N}}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Big{|}\leq\frac{C}{\varepsilon N(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}+\frac{C}{\sqrt{\varepsilon N}(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})^{3/2}}\leq\frac{C}{\sqrt{\varepsilon N}(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})^{3/2}}.$
@@
$\begin{split}\bigg{\|}&\frac{1}{\sqrt{N}}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon,}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}q_{0,d_{1}}^{N}(\varphi,0)X_{d_{1},f_{1}}\Big{(}p_{d_{2}-f_{1}}(0)-\frac{1}{\sqrt{\varepsilon N}}g_{(\mathrm{i}_{2}-\mathrm{i}_{1})}(0)\Big{)}X_{d_{2},f_{2}}\\ &\times\bigg{\{}\prod\limits_{j=3}^{k}p_{d_{j}-f_{j-1}}(0)X_{d_{j},f_{j}}\bigg{\}}q_{f_{k},N}^{N}(0,\psi)\bigg{\|}_{2}^{2}\\ \leq&\frac{C\|\psi\|_{\infty}^{2}}{N}\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\big{(}q_{0,d_{1}}^{N}(\varphi,0)\big{)}^{2}\overline{U}_{N}(f_{1}-d_{1})\frac{1}{\varepsilon N(\mathrm{i}_{2}-\mathrm{i}_{1})^{3}}\overline{U}_{N}(f_{2}-d_{2})\\ &\times\bigg{\{}\prod\limits_{j=3}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\bigg{\}}.\end{split}$
Note that $1/(\varepsilon N(\mathrm{i}_{2}-\mathrm{i}_{1}))\leq C/(d_{2}-f_{1})$ and $\mathrm{i}_{2}-\mathrm{i}_{1}\geq K_{\varepsilon}$ by the “no triple” condition in $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$. By Lemma 5.2, taking $\limsup_{N\to\infty}$, the above term is bounded by $C\|\psi\|_{\infty}^{2}(K_{\varepsilon})^{-2}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$. Since there are at most $2(\log\frac{1}{\varepsilon})^{2}$ kernel replacements, by triangle inequality, the total $L^{2}$ error for kernel replacements is bounded by $C\|\psi\|_{\infty}^{2}(\log\frac{1}{\varepsilon})^{2}(K_{\varepsilon})^{-2}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$. Finally, by summing over all $k$ and $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$, and by the orthogonality of different $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$, the upper bound for kernel replacement is
$\begin{split}&\limsup\limits_{N\to\infty}C\|\psi\|_{\infty}^{2}\big{(}\log\frac{1}{\varepsilon}\big{)}^{2}(K_{\varepsilon})^{-2}{\rm Var}({\mathcal{Z}}_{N}(\varphi,\psi))\\ \leq&\frac{C\big{(}\log\frac{1}{\varepsilon}\big{)}^{2}\|\psi\|_{\infty}^{2}}{K_{\varepsilon}^{2}}\int_{0}^{1}G_{\vartheta}(t){\mathrm{d}}t\int_{0}^{1}g_{s}(\varphi,0)^{2}{\mathrm{d}}s\leq\frac{C_{\vartheta}\big{(}\log\frac{1}{\varepsilon}\big{)}^{2}}{K_{\varepsilon}^{2}}\|\psi\|_{\infty}^{2}\|\varphi\|_{2}^{2}.\end{split}$
@@
Finally we prove (5.12), the approximation to ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}$ by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$. This is just a Riemann sum approximation from $\sum_{\mathrm{a}\sqrt{\varepsilon N}<u\leq(\mathrm{a}+1)\sqrt{\varepsilon N},u\in{\mathbbm{Z}}}\varphi_{N}(u)$ and $\sum_{\mathrm{b}\sqrt{\varepsilon N}<v\leq(\mathrm{b}+1)\sqrt{\varepsilon N},v\in{\mathbbm{Z}}}\psi_{N}(v)$ to $\varphi_{\varepsilon}(\mathrm{a})$ and $\psi_{\varepsilon}(\mathrm{b})$ respectively. ∎
6. Moment estimates for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
In this section, we first bound the second and fourth moments of the coarse-grained disorder $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ (see (2.43)), and then we will use these bounds to bound moments of coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$.
6.1. Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
@@
(6.1) $\sigma_{\varepsilon}^{2}(\vec{\mathrm{i}}):=\lim\limits_{N\to\infty}{\mathbbm{E}}\Big{[}\big{(}\Theta_{N,\varepsilon}(\vec{\mathrm{i}})\big{)}^{2}\Big{]}.$
@@
(6.2) $\sigma_{\varepsilon}^{2}(\vec{\mathrm{i}})\leq\frac{C}{\big{(}\log\frac{1}{\varepsilon}\big{)}^{1+\mathbbm{1}_{\{|\vec{\mathrm{i}}|\geq 2\}}}|\vec{\mathrm{i}}|}.$
@@
The computation is similar to those for Proposition 2.3 and Lemma 5.2. When $|\vec{\mathrm{i}}|\leq 2$, $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ is an average of $X_{d,f}$ (see (2.40)),which has the same structure as ${\mathcal{Z}}_{N}(\varphi,\psi)$ with $\psi(y)=\mathbbm{1}_{0}(y)$. When $|\vec{\mathrm{i}}|\geq 3$, we take advantage of the fact that $p_{d^{\prime}-f}(0)$ in (2.43) is uniformly bounded below by $C(|\vec{\mathrm{i}}|\varepsilon N)^{-1/2}$. ∎
@@
The strategy is the same as in Lemma 6.1 and the difference is that we apply Theorem 4.1. We split the case $|\vec{\mathrm{i}}|\leq 2$ and $|\vec{\mathrm{i}}|\geq 3$. Note that in the latter case, we can decouple the fourth moments of $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ in (2.43) by the independence.∎
6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
@@
$\begin{split}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta):=g_{1}(\varphi,\psi)+&\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta(\vec{\mathrm{i}}_{1})\times\bigg{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta(\vec{\mathrm{i}}_{j})\bigg{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}).\end{split}$
We have the following fourth moment bound for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$, which is an analogue to Theorem 4.1.
@@
where $\varphi_{\varepsilon}$ and $w_{\varepsilon}$ are defined by (2.47).
@@
We introduce $w$ since in Section 7 we will consider $\psi$ with unbounded support.

### v3

@@ 5.1 Approximating ${\mathcal{Z}}_{N}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$
To bound (5.10), a fundamental tool is still the renewal property (5.8). For the cases $\mathrm{i}_{1}<K_{\varepsilon}$ and $\mathrm{i}_{k}>\frac{1}{\varepsilon}-K_{\varepsilon}$, since each ${\mathcal{T}}_{\varepsilon}(\mathrm{i})$ is of length $\varepsilon$, we have an integral on $0,K_{\varepsilon}\varepsilon]$, whose length tends to 0 by the choice of $K_{\varepsilon}$ in (2.38). For the case that there exists consecutive $\mathrm{i}-\mathrm{i}^{\prime}<K_{\varepsilon}$ and $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}<K_{\varepsilon}$, we need to consider separately the cases (1) $\mathrm{i}^{\prime}-\mathrm{i}\geq 2$ and $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}\geq 2$; (2) $\mathrm{i}^{\prime}-\mathrm{i}=1$ or $\mathrm{i}^{\prime\prime}-\mathrm{i}^{\prime}=1$. The case (1) is relatively simple, since we can bound the transition kernels between mesoscopic intervals from below uniformly. The case (2) needs a delicate computation, the readers are recommended to refer to [15, Section 5.1] for details. ∎
5.2 Approximating ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}$ by ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}$ and $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
@@
$\limsup\limits_{N\to\infty}$ $\left\|\left({\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(no triple)}}-{\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}\right)(\varphi,\psi)\right\|_{2}^{2}\leq C\Big(\frac{1}{\log\frac{1}{\varepsilon}}+\frac{(\log\frac{1}{\varepsilon})^{2}}{K_{\varepsilon}^{\kappa}}\Big)\|\varphi\|_{2}^{2}\|\psi\|_{\infty}^{2},$ (5.11)
$\lim\limits_{N\to\infty}$ $\left\|\left({\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}-\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)\right)(\varphi,\psi)\right\|_{2}^{2}=0.$ (5.12)
@@
Step 1. To expand ${\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$, we bind $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ together via a transition probability $p_{d^{\prime}-f}(0)$, where $d,f\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i})$ and $d^{\prime},f^{\prime}\in{\mathcal{T}}_{\varepsilon,N}(\mathrm{i}^{\prime})$ with $\mathrm{i}^{\prime}-\mathrm{i}<K_{\varepsilon}$. Since each $\vec{\mathrm{i}}=(\mathrm{i},\mathrm{i}^{\prime})$ can at most refer to 2 mesoscopic intervals. By expanding ${\mathcal{A}}_{\varepsilon}^{\text{\rm(no triple)}}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$, the terms refer to indices ${\mathbf{I}}=(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$ with $(\log\frac{1}{\varepsilon})^{2}\leq k\leq 2(\log\frac{1}{\varepsilon})^{2}$ will be added, and the error has been controlled by (5.3) and (5.10).
Step 2. Then we replace the transition kernels by the heat kernels. Note that each time we bind two $X_{d_{j},f_{j}}$ and $X_{d^{\prime}_{j},f^{\prime}_{j}}$ together to get a $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})$, we have an extra $\sqrt{\varepsilon N}$. For the kernel connecting $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j-1})$ and $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})$, we have an approximation $\sqrt{\varepsilon N}p_{d_{j}-f^{\prime}_{j-1}}(0)\approx g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)$. We need to show that such kernel replacement is controllable. By Theorem B.1, for fixed $\varepsilon>0$ and $N$ large enough,
$\begin{split}\Big|p_{d_{j}-f^{\prime}_{j-1}}(0)-&\frac{1}{\sqrt{\varepsilon N}}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Big|\leq\frac{C}{(d_{j}-f^{\prime}_{j-1})^{(1+\kappa)/2}}+\Big|g_{d_{j}-f^{\prime}_{j-1}}(0)-\frac{1}{\sqrt{\varepsilon N}}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Big|\\ \leq&\frac{C}{(d_{j}-f^{\prime}_{j-1})^{(1+\kappa)/2}}+\frac{C}{\sqrt{\varepsilon N}(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})^{3/2}}\leq\frac{C}{\sqrt{\varepsilon N}(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})^{(1+\kappa)/2}}.\end{split}$ (5.13)
@@
$\begin{split}\bigg\|&\frac{1}{\sqrt{N}}\!\!\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon,}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\!\!q_{0,d_{1}}^{N}(\varphi,0)X_{d_{1},f_{1}}\Big(p_{d_{2}-f_{1}}(0)-\frac{1}{\sqrt{\varepsilon N}}g_{(\mathrm{i}_{2}-\mathrm{i}_{1})}(0)\Big)X_{d_{2},f_{2}}\\ &\times\bigg\{\prod\limits_{j=3}^{k}p_{d_{j}-f_{j-1}}(0)X_{d_{j},f_{j}}\bigg\}q_{f_{k},N}^{N}(0,\psi)\bigg\|_{2}^{2}\\ \leq&\frac{C\|\psi\|_{\infty}^{2}}{N}\!\!\!\!\!\!\sum\limits_{d_{1}\leq f_{1}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{1}),\cdots,d_{k}\leq f_{k}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}_{k})}\!\!\!\!\!\!\!\big(q_{0,d_{1}}^{N}(\varphi,0)\big)^{2}\overline{U}_{N}(f_{1}-d_{1})\frac{1}{\varepsilon N(\mathrm{i}_{2}-\mathrm{i}_{1})^{1+\kappa}}\overline{U}_{N}(f_{2}-d_{2})\\ &\times\bigg\{\prod\limits_{j=3}^{k}u(d_{j}-f_{j-1})\overline{U}_{N}(f_{j}-d_{j})\bigg\}.\end{split}$
Note that $1/(\varepsilon N(\mathrm{i}_{2}-\mathrm{i}_{1}))\leq C/(d_{2}-f_{1})$ and $\mathrm{i}_{2}-\mathrm{i}_{1}\geq K_{\varepsilon}$ by the “no triple” condition in $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$. By Lemma 5.2, taking $\limsup_{N\to\infty}$, the above term is bounded by $C\|\psi\|_{\infty}^{2}(K_{\varepsilon})^{-\kappa}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$. Since there are at most $2(\log\frac{1}{\varepsilon})^{2}$ kernel replacements, by triangular inequality, the total $L^{2}$ error for kernel replacements is bounded by $C\|\psi\|_{\infty}^{2}$ timing $(\log\frac{1}{\varepsilon})^{2}(K_{\varepsilon})^{-\kappa}{\mathcal{I}}_{\varepsilon}^{\varphi}(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$. Finally, by summing over all $k$ and $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$, and by the orthogonality of different $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$, the upper bound for kernel replacements is
$\begin{split}&\limsup\limits_{N\to\infty}C\|\psi\|_{\infty}^{2}\big(\log\frac{1}{\varepsilon}\big)^{2}(K_{\varepsilon})^{-\kappa}{\rm Var}({\mathcal{Z}}_{N}(\varphi,\psi))\\ \leq&\frac{C\big(\log\frac{1}{\varepsilon}\big)^{2}\|\psi\|_{\infty}^{2}}{K_{\varepsilon}^{\kappa}}\int_{0}^{1}G_{\vartheta}(t){\mathrm{d}}t\int_{0}^{1}g_{s}(\varphi,0)^{2}{\mathrm{d}}s\leq\frac{C_{\vartheta}\big(\log\frac{1}{\varepsilon}\big)^{2}}{K_{\varepsilon}^{\kappa}}\|\psi\|_{\infty}^{2}\|\varphi\|_{2}^{2}.\end{split}$
@@
Finally we prove (5.12), the approximation to ${\mathcal{Z}}_{N,\varepsilon}^{\text{\rm(cg)}}$ by $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$. This is just a Riemann sum approximation from $\sum\limits_{\mathrm{a}\sqrt{\varepsilon N}<u\leq(\mathrm{a}+1)\sqrt{\varepsilon N},u\in{\mathbbm{Z}}}\varphi_{N}(u)$ and $\sum\limits_{\mathrm{b}\sqrt{\varepsilon N}<v\leq(\mathrm{b}+1)\sqrt{\varepsilon N},v\in{\mathbbm{Z}}}\psi_{N}(v)$ to $\varphi_{\varepsilon}(\mathrm{a})$ and $\psi_{\varepsilon}(\mathrm{b})$ respectively. ∎
6 Moment estimates for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$
In this section, we first bound the second and fourth moments of the coarse-grained disorder $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ (see (2.43)), and then we will use these bounds to bound the moments of the coarse-grained model $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$.
6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
@@
$\sigma_{\varepsilon}^{2}(\vec{\mathrm{i}}):=\lim\limits_{N\to\infty}{\mathbbm{E}}\Big[\big(\Theta_{N,\varepsilon}(\vec{\mathrm{i}})\big)^{2}\Big].$ (6.1)
@@
$\sigma_{\varepsilon}^{2}(\vec{\mathrm{i}})\leq\frac{C}{\big(\log\frac{1}{\varepsilon}\big)^{1+\mathbbm{1}_{\{|\vec{\mathrm{i}}|\geq 2\}}}|\vec{\mathrm{i}}|}.$ (6.2)
@@
The computation is similar to those for Proposition 2.3 and Lemma 5.2. When $|\vec{\mathrm{i}}|\leq 2$, $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ is an average of $X_{d,f}$ (see (2.40)),which has the same structure as ${\mathcal{Z}}_{N}(\varphi,\psi)$ with $\varphi(x)=\mathbbm{1}_{0}(x)$ and $\psi(y)=\mathbbm{1}_{0}(y)$. When $|\vec{\mathrm{i}}|\geq 3$, we take advantage of the fact that $p_{d^{\prime}-f}(0)$ in (2.43) is uniformly bounded below by $C(|\vec{\mathrm{i}}|\varepsilon N)^{-1/2}$ and decouple $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ by the independence of $\zeta_{n}$. ∎
@@
The strategy is the same as in Lemma 6.1 and the difference is that we apply Theorem 4.1. We split the cases $|\vec{\mathrm{i}}|\leq 2$ and $|\vec{\mathrm{i}}|\geq 3$. Note that in the former case, the length of polymer chains is $|\vec{\mathrm{i}}|\varepsilon N$ and we can apply Theorem 4.1 with $\varphi(x)=\mathbbm{1}_{0}(x)$ and $\psi(y)=\mathbbm{1}_{0}(y)$ (see Remark 4.2). In the latter case, as in the proof of Lemma 6.1, we can decouple the fourth moments of $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ in (2.43) by the independence.∎
6.2 Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$
@@
$\begin{split}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)&:=g_{1}(\varphi,\psi)+\sqrt{\varepsilon}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta(\vec{\mathrm{i}}_{1})\times\bigg\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta(\vec{\mathrm{i}}_{j})\bigg\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b}).\end{split}$
We have the following fourth moment bound for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)=\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi|\Theta)$, which is an analogue to Theorem 4.1.
@@
where $\varphi_{\varepsilon}$ and $w_{\varepsilon}$ are defined through (2.47).
@@
We introduce $w$ since in Section 7 we will consider $\psi$ with unbounded support. A typical choice of $w(x)$ is $e^{-|x|}$.